Presidio, Inc., produces one model of a mountain bike. Partial information for the company follows:

Required:
1. Complete Presidio’s cost data table.
2. Calculate Presidio’s contribution margin ratio and its total contribution margin at each sales level indicated in the cost data table assuming the company sells each bike for $630.
3. Calculate net operating income (loss) at each of the sales levels assuming a sales price of $630.

Complete Presidio’s cost data table. (Round your Cost per Unit answers to 2 decimal places.)

places (i.e. .1234 should be entered as 12.34%.))

Calculate net operating income (loss) at each of the sales levels assuming a sales price of $630. (Round your answers to the nearest whole dollar amount.)

3.
720 Units 750 Units 1,248 Units
Net Operating Income
2.
720 Units 750 Units 1,248 Units
Total Contribution Margin
Contribution Margin Ratio % % %
1.
Bikes Produced and Sold 720 Units 750 Units 1,248 Units
Total costs
Variable costs $180,000
Fixed costs per year
Total costs $180,000 $0 $0
Cost per unit
Variable cost per unit
Fixed cost per unit
Total cost per unit $0.00 $544.00 $0.00
Answer #1
1.) Amount $
Bikes Produced and Sold 720 Units 750 Units 1,248 Units
Total costs
Variable costs               180,000              187,500               312,000
Fixed costs per year               220,500              220,500               220,500
Total costs               400,500              408,000               532,500
Cost per unit
Variable cost per unit                  250.00                 250.00                  250.00
Fixed cost per unit                  306.25                 294.00                  176.68
Total cost per unit                  556.25                 544.00                  426.68
Working:
Bikes Produced and Sold 720 Units 750 Units 1,248 Units
Total costs
Variable costs               180,000 =750*250 =1248*250
Fixed costs per year               220,500 =408000-187500               2,20,500
Total costs =180000+220500 =750*544 =220500+312000
Cost per unit
Variable cost per unit =180000/720 =187500/750 =312000/1248
Fixed cost per unit =220500/720 =220500/750 =220500/1248
Total cost per unit =250+306.25                 544.00 =250+176.68
2.) 720 Units 750 Units 1,248 Units
Total Contribution Margin $               273,600              285,000               474,240
(Selling price per unit - Variable Cost per unit ) x unit solds
Contribution Margin Ratio 60.32% 60.32% 60.32%
( Contribution margin / Sales )
Working
720 Units 750 Units 1,248 Units
Total Contribution Margin =(630-250)*720 =(630-250)*750 =(630-250)*1248
Contribution Margin Ratio =273600/(720*630) =285000/(750*630) =474240/(1248*630)
3.) 720 Units 750 Units 1,248 Units
Net Operating income                       53,100                      64,500                   253,740
( Contribution margin - Fixed Costs )
Working
720 Units 750 Units 1,248 Units
Net Operating income =273600-220500 =285000-220500 =474240-220500